quickstudent:

The sine of 12 degrees is equal to 0.208, rounded to the nearest thousandth. What is the measure of an angle whose cosine is also equal to 0.208? 12° 68° 78° 102°

Hero:

Hint: \(\cos(\theta) = \sin(90 - \theta)\)

quickstudent:

78

Hero:

Correct

Hero:

Also, check your previous question.